Previous Problem List (I point) Consider the following frequency distribution. Next Class Frequency 8-19 19-30 7 30-41 8 41-52 6 2 What is the midpoint of the first class? To the nearest hundredth, what is the relative frequency of the third class? 24 Note: You can earn partial credit on this problem. (Give a decimal number not a percentage,) You have attempted this problem 0 times. You have unlimited attempts remaining.
Answer #1

Please don't hesitate to give a "thumbs up " for the answer, in case you're satisfied with it.

Total frequency is 4+7+8+6 = 25.

Mid point of 1st class = (8+19)/2 = 13.5

Relative frequency of 3rd class is it' frequency divided by the total frequency = 8/(25) = .32